# Polynomial-like dynamics of analytic maps

### v1

@@ 3.1. The case of connected $K$
The latter condition implies that $L$ must be homotopic to a point in the domain $\hat{\mathbb{C}}\setminus K$. On the other hand, as the curve $l$ lies in a simply connected domain $Q$ which is disjoint with some $Q_{j}$, the curve $L$ must lie in one of the components of $g^{-n}(Q)$, which, in turn, are disjoint with all components of $g^{-n}(Q_{j})$. As each of the latter components contains points of $K$, the curve $L$ cannot enclose $\infty$ (in $\hat{\mathbb{C}}$), hence, $L$ is homotopic to a point in the subdomain $\mathbb{C}\setminus K$ of $\mathbb{C}$. Since $K$ is completely invariant, then $l$ is homotopic to a point in $\mathbb{C}\setminus K$ too, hence, its lift $\hat{l}$ must be a closed curve, a contradiction. This proves Claim 1.
Thus, for every $n\in\mathbb{N}$, $\hat{\psi}_{n}:\Omega\to\mathbb{D}$ is a univalent functions such that $\hat{\psi}_{n}|_{\hat{Z}}=\hat{\psi}^{n}$. Let $\hat{S}:\hat{Z}\to\mathbb{C}$ be the Konigs linearization function of $\hat{\psi}$ at its attracting fixed point $\hat{a}$, so that $\hat{S}(\hat{a})=0$, $\hat{S}^{\prime}(\hat{a})\neq 0$ and $\hat{S}(w)=\lambda\hat{S}(\hat{\psi}(w))$ for $w\in\hat{Z}$. Let $\hat{S}=S\circ\pi$ on $\hat{Z}$.
Our goal is to prove that
  1. (i)
    $\hat{S}$ extends to a holomorphic function to $\Omega$ (denoted also by $\hat{S}$), and
  2. (ii)
    $\hat{S}(w)\to\infty$ as $w\to F_{0}=\partial\Omega\cap\partial\mathbb{D}$.
As the boundary of the simply connected domain $\Omega=W\cup T\subset\mathbb{D}$ is a rectifiable curve and the length of $F_{0}=\Omega\cap\partial\mathbb{D}=\partial W\cap\partial\mathbb{D}$ is positive, (i)-(ii) would be in a contradiction with Privalov’s theorem.
Proof of (i): choose an increasing sequence $W_{i}$, $i\geq 1$, of subdomains of $W$ along with an increasing sequence of "tubes" $T_{i}$, such that, for $\Omega_{i}=T_{i}\cup W_{i}$, $\overline{\Omega_{i}}\subset\Omega$ for all $i$ and $\cup_{i}\Omega_{i}=\Omega$. As $(\hat{\psi}_{n})^{\prime}(\hat{a})=\lambda^{-n}\to 0$ as $n\to\infty$ and $\hat{\psi}_{n}(\hat{a})=\hat{a}$, by Koebe’s distortion theorem applied to maps $\psi_{n}:\Omega\to\mathbb{C}$, for any fixed $i$, $\hat{\psi}_{n}(W_{i})\to\hat{a}$ as $n\to\infty$. Choose some $n_{i}$ such that $\hat{\psi}_{n_{i}}(\Omega_{i})\subset\hat{Z}$ and then extend the function $\hat{S}:\hat{Z}\to\mathbb{C}$ to some function $\hat{S}_{i}:\Omega_{i}\to\mathbb{C}$ as follows: $\hat{S}_{i}(w)=\lambda^{n}\hat{S}(\hat{\psi}_{n_{i}}(w))$ for $w\in\Omega_{i}$. It is easy to check that $\hat{S}_{i}$ is well-defined and holomorphic in $\Omega_{i}$ and $\hat{S}_{i}=\hat{S}$ on $\hat{Z}$. Then, by the Uniqueness theorem, $\hat{S}_{i}=\hat{S}_{i+1}$ on $\Omega_{i}$. That proves (i).
Thus $\hat{S}:\Omega\to\mathbb{C}$ such that $S|_{\Omega_{i}}=S_{i}$ defines a holomorphic extension of $\hat{S}:\hat{Z}\to\mathbb{C}$.
Proof of (ii): let $w_{k}\in\Omega$, $k=1,2,$..., be any sequence that tends to $F_{0}=\partial\Omega\cap\partial\mathbb{D}$. As $F_{0}=\partial W\cap\partial\mathbb{D}$, then all but finitely many of $w_{k}$ are in $W$. Let $w_{k}\in W_{i_{k}}$ and $m_{k}:=n_{i_{k}}$ so that $\hat{\psi}_{m_{k}}(W_{i_{k}})\subset\hat{Z}$ and $\hat{S}(w_{k})=\lambda^{m_{k}}\hat{S}(\hat{\psi}_{m_{k}}(w_{k}))$. Let
[…]
in other words,
[…]
Claim 2. $e_{k}\leq m_{k}$ for all $k$ large enough.
[…]
Claim 3. $m_{k}-e_{k}\to\infty$ as $k\to\infty$.
[…]
(3.1) $g^{-(m_{k}-e_{k})}|_{loc}(z_{k})\in Z.$
By Corollary B.1(6) (applied to $X=K$), as $w_{k}\to\partial\mathbb{D}$ inside of $W$, then $z_{k}=\pi(w_{k})\to K$. But the open set $Z$ is of a positive distance away from the compact $K$ and $g^{-n}(K)=K$ for all $n$. Then (3.1) can hold only if $m_{k}-e_{k}\to+\infty$.
End of proof of (ii):
$\hat{S}(w_{k})=\lambda^{m_{k}}\hat{S}(\hat{\psi}_{m_{k}}(w_{k}))\in$
[…]
@@
Choose a Jordan domain $V_{0}\subset U$ such that $K\subset V_{0}$, therefore, $P_{g}\subset V_{0}$ as well. By Theorem 1.2, the alternative (B) of Proposition 2.2 holds, which implies that the conclusion of Theorem 1.2 holds.
5. Proof of Corollary 1.1
Let $Y$ be any non-empty completely invariant full compact set.
@@
(2) If $P_{g}\subset Y$, then Theorem 1.1 implies that $\partial Y$ is the filled Julia set of a polynomial-like restriction to a neighborhood $P_{g}$. This implies that $Y$ must coincide with the limit set of $\{g^{-n}(z)\}_{n>0}$, where $z$ can be taken to be e.g. a critical point of $g:U^{\prime}\to U$. Hence, $Y$ is unique in this case, i.e., $K_{*}=Y$. To prove the second claim, let $x\in U$ be such that $O(x)$ is well-defined but $\omega(x)\nsubseteq\partial U$. In other words, there exist a sequence $n_{i}\to\infty$ a compact $E\subset U$ so that $g^{n_{i}}(x)\in E$, for all $i$. Choose a Jordan domain $V$ such that $P_{g}\cup E\subset V\subset\overline{V}\subset U$. As $P_{g}\subset K_{*}$, by Theorem 1.2 along with Proposition 2.2, the alternative (B) of Proposition 2.2 must hold, i.e., for some $n\in\mathbb{N}$, $V_{-n}:=g^{-n}(V_{0})$ is contained in $V$ with its closure. By (B), then $K_{*}=K_{G}:=\cap_{k\geq 0}g^{-kn}(V)$. On the other hand, write $n_{i}=l_{i}n+r_{i}$, with integers $l_{i},r_{i}$ such that $l_{i}\to\infty$ and $0\leq r_{i}<n$. Passing to a subsequence, one can assume that $r_{i}=r$ for all $i$. Then $g^{l_{i}n}(g^{r}(x))\in E\subset V$, for all $i$, hence, $g^{r}(x)\in\cap_{i\geq 0}g^{-l_{i}n}(V)\subset\cap_{k\geq 0}g^{-kn}(V)$, where the latter equality follows from the fact that the sequence of domains $\{g^{-kn}(V)\}_{k\geq 0}$ is decreasing. Thus $g^{r}(x)\in K_{*}$ so that $x\in g^{-r}(K_{*})=K_{*}$.
(3) Assume that $P_{g}\nsubseteq Y$. Let $X=P_{g}\cup Y$. Then $X$ is a forward invariant compact subset of $U$ which contains $P_{g}\cup Y$. Define $X_{*}$ as in Proposition 2.5. Then $Y\subset X_{*}$. On the other hand, by Proposition 2.5, either $P_{g}\subset X_{*}$ and then, by part (1) of the proof, $Y=X_{*}\supset P_{g}$, a contradiction, or $\partial X_{*}\subset P_{g}$. The latter implies that $K_{*}=X_{*}=\{z:GO(z)\subset\hat{P}_{g}\}\supset Y$. The conclusion follows.
Now, let $Z$ be any non-empty completely invariant compact set and let, similar to above, $X=Z\cup P_{g}$. Then $X$ is forward invariant. Define $X_{*}$ as in Proposition 2.5. By Proposition 2.5 and by the proven part, $K_{*}=X_{*}$. On the other hand, $Z\subset X_{*}$.
The final claim involving the set $R_{g}$ follows directly from Corollary 2.3 in cases (1) and (3), and from the Straightening theorem coupled with the fact that repelling cycles are dense in the polynomial Julia set in case (2).
@@
The proof follows almost literally the one of Theorem 1.1 for disconnected $K$, where very minor changed are required and only in Proposition 2.2 and its proof. First, one should restrict it only to Jordan curves $\Gamma_{0}$ which are close enough to $\partial U$. Curves $\Gamma_{n}$ turn into multi-curves. In the alternative (C), $\overline{V_{0}}$ should be a subset of a component of $V_{-n}$. In the part of the proof of Proposition 2.2 where (A) holds, the normality argument in the case $P_{g}=\emptyset$ can be easily amended; in the part where (B) holds, $V$ should be a neighborhood of $K$ consisting of a finitely many simply connected components and the Second proof of the construction of $V$ such that $g^{-1}(V)$ is compactly in $V$ should be chosen.
@@ Appendix A Entropy of locally analytic maps
(2) If, additionally, $f:W^{\prime}\to W$ is a proper map of degree $N$ where $W$ is simply connected and $W^{\prime}$ consists of a finitely many simply connected components, then the conclusion holds unless $W^{\prime}$ is (simply) connected and $X$ is a singleton $\{c\}$ where $c$ is a fixed critical point of $f$ of multiplicity $N$.
@@ Appendix B Uniformization
For a Borel subset $F\subset\partial\mathbb{D}$, let $|F|$ be its Lebesgue measure (length). We have:
@@
Now we repeat partly a construction as in the proof of Privalov’s theorem, e.g. [8]. Fix an angle $\beta\in(0,\pi/2)$ and given $a=\exp(i\theta)\in\partial\mathbb{D}$ and $n\in\mathbb{N}$ define
$H_{n}(a)=\inf\{|B(w)|:1-1/n<|w|<1,|\arg(w-a)|<\beta\}.$

### v2

@@ 4.1. The case of connected $K$
The latter condition implies that $L$ must be homotopic to a point in the domain $\hat{\mathbb{C}}\setminus K$. On the other hand, as the curve $l$ lies in a simply connected domain $Q$ which is disjoint with some $Q_{j}$, the curve $L$ must lie in one of the components of $g^{-n_{i}}(Q)$, which, in turn, are disjoint with all components of $g^{-n_{i}}(Q_{j})$. As each of the latter components contains points of $K$, the curve $L$ cannot enclose $\infty$ (in $\hat{\mathbb{C}}$), hence, $L$ is homotopic to a point in the subdomain $\mathbb{C}\setminus K$ of $\mathbb{C}$. Since $K$ is completely invariant, then $l$ is homotopic to a point in $\mathbb{C}\setminus K$ too, hence, its lift $\hat{l}$ must be a closed curve, a contradiction. This proves Claim 1.
Let $F=\partial W\cap\partial\mathbb{D}$. Thus, we have:
  1. ($\hat{1}$)
    for every big $i$, $\hat{\psi}_{i}:\Omega\to\mathbb{D}$ is a univalent continuation of $\hat{\psi}_{i}:\hat{Z}\to\hat{\mathbb{Z}}$,
  2. ($\hat{2}$)
    $\hat{\psi}_{i}(\hat{a}_{i})=\hat{a}_{i}$, $\hat{a}_{i}\to\hat{a}\in\Omega$ and $\hat{\psi}_{i}^{\prime}(\hat{a}_{i})=\psi_{i}^{\prime}(a_{i})\to 0$,
  3. ($\hat{3}$)
    as $|F|>0$ and $K$ is completely invariant, for every big $i$ and almost every $w_{0}\in F$, $\hat{\psi}_{i}(w)$ has a an angular limit $\hat{\psi}(w_{0})$ as $w\to w_{0}$ in $\Omega$ and $\hat{\psi}_{i}(w_{0})\in\partial\mathbb{D}$.
We claim that ($\hat{1}$)-($\hat{3}$) are incompatible.
Indeed, let $R:\mathbb{D}\to\Omega$ be a Riemann map such that $R(0)=\hat{a}$. Let $f_{i}=\hat{\psi}_{i}\circ R$, a univalent map in $\mathbb{D}$, for each $i$. As $\partial\Omega$ is a rectifiable Jordan curve, $R^{\prime}\in H^{1}(\mathbb{D})$. Hence, $E:=R^{-1}(F)$ is a measurable subset of $S^{1}$ of a positive 1-dim measure $|E|>0$.
Let $f_{i}=\hat{\psi}_{i}\circ R$, a univalent map of $\mathbb{D}$ onto a bounded domain in $\mathbb{D}$, for each $i$. Since $\hat{\psi}_{i}:\Omega\to\mathbb{D}$ are univalent and form a normal family, by ($\hat{2}$), $f_{i}(0)=\hat{\psi}_{i}(\hat{a})\to\hat{a}$ where $\hat{a}\in\Omega\subset\mathbb{D}$, and $f_{i}^{\prime}(0)=\hat{\psi}_{i}^{\prime}(\hat{a})R^{\prime}(0)\to 0$. By ($\hat{3}$), $f_{i}(E)\subset S^{1}$, for all $i$. Therefore, one can choose $i_{*}$ such that
$|E|\sqrt{\frac{(1-|f_{i_{*}}(0)|)^{3}}{(1+|f_{i_{*}}(0)|)|f_{i_{*}}^{\prime}(0)|}}>2\pi,$
in a contradiction with Corollary C.3 applied to $f=f_{i_{*}}$.
@@
It now follows directly from Theorem 1.2 along with Proposition 3.2, where the alternative (A) must occur.
6. Proof of Corollary 1.5 and its Complement
Proof of (b). By Proposition 3.2, for every Jordan curve $\Gamma$ in $U_{0}$ which encloses $P_{g}$ either (A) or (C) must hold. If (C) holds for a sequence of such $\Gamma_{0}$ that tends to $\partial U$, then, by (C), $g:U^{\prime}\to U$ is trivial, which is ruled out by a condition. Therefore, for each Jordan curve $\Gamma$ which surrounds $P_{g}$ and is close enough to $\partial U_{0}$ the alternative (A) holds, that is, $\Gamma_{0}$ must intersect $R^{\prime}$. Now, consider a connected component $C$ of $R_{g}\cup\partial U_{0}$ which contains $\partial U_{0}$ It is enough to prove that it also contains points of $U_{0}$. Assume the contrary, i.e., $C=\partial U$. Fix some Jordan curve $\Gamma^{\prime}$ as above (i.e., $\Gamma^{\prime}$ surrounds $P_{g}$ and close to $\partial U_{0}$) and let $C^{\prime}$ be a connected component of $R_{g}\cup\Gamma^{\prime}$ that contains $\Gamma^{\prime}$. Let $\Omega$ be the unbounded component of $\mathbb{C}\setminus C^{\prime}$ and $D=(\Omega\cap U_{0})\setminus R^{\prime}$. By the assumption, $D$ is an open connected set and $\partial U_{0}$ is a component of the boundary $\partial D$ while $\partial\Omega$ is another component of $\partial D$. Therefore, there exists a curve $\Gamma^{\prime\prime}$ in $D$ which surrounds $\partial\Omega$, hence, $P_{g}$ as well, and $\Gamma^{\prime\prime}$ is close to $\partial U_{0}$. Then $\Gamma^{\prime\prime}$ must intersect $R^{\prime}$, a contradiction because $\Gamma^{\prime\prime}\subset D$ and $D$ is disjoint with $R^{\prime}$.
Proof of (a). Item (1): by Theorem 1.1, $K$ is the filled Julia set of a PL restriction $g:g_{0}^{-1}(W)\to W$. By the Straightening theorem and the Fatou-Julia theory, $\partial K$ is equal to the limit set of repelling periodic points of $g:K\to K$. On the other hand, by Lemma 3.1 applied to $K$, $R_{n}\cap K$ coincides with the set $\{z\in K:g^{n}(z)=z,|(g^{n})^{\prime}(z)|>1\}$. Hence, $R^{\prime}\cap K=\partial K$, i.e., item (1). Item (2): assume, by a contradiction, that $x\in W\setminus K$ and $g^{n}(x)=x$ for some $n\in\mathbb{N}$ and $x\in U_{-n}\cap U_{0}$. On the one hand, $x\in g_{0}^{-n}(x)$. On the other hand, $\overline{g_{0}^{-1}(W)}\subset W$. Let $A_{k}=g_{0}^{-k}(W)\setminus\overline{g_{0}^{-k-1}(W)}$, $k=0,1,2,$.... Then $\{A_{n}\}_{n\geq 0}$ is a partition of $W\setminus K$. If now $x\in A_{j}$, for some $j\geq 0$, then $x\in g_{0}^{-n}(x)\in A_{j+n}$ while $A_{j}\cap A_{j+n}=\emptyset$, a contradiction. Item (3) follows directly from Theorem 1.2.
@@
Proof of the Complement. After the proven Corollary and Remark 1.6, we only need to show the claim about iterates of $x$. Let $x\in V^{\prime}\setminus K$ be such that $O(x)$ is well-defined but $\omega(x)\nsubseteq\partial V$. In other words, there exist a sequence $n_{i}\to\infty$ a compact $E\subset V$ so that $g^{n_{i}}(x)\in E$, for all $i$. Choose a Jordan domain $V_{0}$ such that $P_{g}\cup E\subset V_{0}\subset\overline{V_{0}}\subset V$. As $P_{g}\subset K$, by Theorem 1.2 along with Proposition 3.2, the alternative (B) of Proposition 3.2 must hold, i.e., for some $n\in\mathbb{N}$, $V_{-n}:=g^{-n}(V_{0})$ is contained in $V_{0}$ with its closure. By (B), then $K=K_{G}:=\cap_{k\geq 0}g^{-kn}(V)$. On the other hand, write $n_{i}=l_{i}n+r_{i}$, with integers $l_{i},r_{i}$ such that $l_{i}\to\infty$ and $0\leq r_{i}<n$. Passing to a subsequence, one can assume that $r_{i}=r$ for all $i$. Then $g^{l_{i}n}(g^{r}(x))\in E\subset V_{0}$, for all $i$, hence, $g^{r}(x)\in\cap_{i\geq 0}g^{-l_{i}n}(V_{0})\subset\cap_{k\geq 0}g^{-kn}(V_{0})$, where the latter equality follows from the fact that the sequence of domains $\{g^{-kn}(V_{0})\}_{k\geq 0}$ is decreasing. Thus $g^{r}(x)\in K$ so that $x\in g^{-r}(K)=K$, a contradiction.
@@
The proof follows almost literally the one of Theorem 1.1 for disconnected $K$, where very minor changed are required and only in Proposition 3.2 and its proof. Curves $\Gamma_{n}$ turn into multi-curves. In the alternative (C), $\overline{V_{0}}$ should be a subset of a component of $V_{-n}$. In the part of the proof of Proposition 3.2 where (A) holds, the normality argument in the case $P_{g}=\emptyset$ can be easily amended; in the part where (B) holds, $V$ should be a neighborhood of $K$ consisting of a finitely many simply connected components and the Second proof of the construction of $V$ such that $g^{-1}(V)$ is compactly in $V$ should be chosen.
@@ Appendix A Entropy of locally analytic maps
(2) If, additionally, $f:W^{\prime}\to W$ is a proper map of degree $N$ where $W$ is simply connected, then the conclusion holds unless $W^{\prime}$ is (simply) connected and $X$ is a singleton $\{c\}$ where $c$ is a fixed critical point of $f$ of multiplicity $N$.
@@ Appendix B Uniformization
Definition B.1.
For an angle $\beta\in(0,\pi/2)$, $r\in(0,\sin\beta)$, and a point $a\in\partial\mathbb{D}$, let $\Delta_{a}(\beta,r)$ be a connected component of the set $\{z:r<|z|<1,|\arg(z-a)|<\beta\}$ which contains $a$ in its boundary.
In other words, $\Delta_{a}(\beta,r)$ is an open "triangle" at a vertex $a$ bounded by an arc of the circle $|z|<r$ and segments of two rays through $a$ which form the angle $\beta$ with the radius $[0,a]$ of the unit circle.
Recall that $|F|$ denotes the Lebesgue 1-dim measure (length) of a measurable subset $F\subset\partial\mathbb{D}$
We have:
@@
Now we repeat partly a Lusin-Privalov construction (as in the proof of Privalov’s theorem, e.g. [17]). Fix an angle $\beta\in(0,\pi/2)$ and given $a\in\partial\mathbb{D}$ and $n\in\mathbb{N}$, such that $1-1/n<\sin\beta$ define
$H_{n}(a)=\inf\{|B(w)|:w\in\Delta_{a}(\beta,1-1/n)\}.$
@@
Appendix C Boundary distortion of sets under univalent self-maps of a disk
Let $f:\mathbb{D}\to\mathbb{D}$ be a univalent map of the unit disk into itself. The following are known theorems, see [11].
  • For a.e. $\zeta\in S^{1}=\partial\mathbb{D}$, $f(z)$ has an angular limit $f(\zeta)$.
  • $f$ has a finite angular derivative at $\zeta\in S^{1}$ if and only if the function $f^{\prime}(z)$ has an angular limit $f^{\prime}(\zeta)$ at $\zeta$.
  • If, for some $\zeta\in S^{1}$, $f(\zeta)\in S^{1}$, then (the angular derivative) $f^{\prime}(\zeta)$ exists and $1\leq\frac{\zeta}{f(\zeta)}f^{\prime}(\zeta)\leq\infty$ (the Wolff-Denjoy Theorem).
  • If $E\subset S^{1}$ is measurable and $f(E)\subset S^{1}$, then $f^{\prime}(\zeta)$ is finite for a.e. $\zeta\in E$ (a corollary from the McMillan Twist Theorem).
In the next lemma, the inequality (C.1) of part 1 is proved in [10] [^1] Although not explicitly stated in [10], the inequality (C.1) is attained as the limit case, as $\to 1, zz\in\mathbb{R}$, in the inequality at the top of p.259 assuming w.l.o.g. that $\zeta=1$., [15], [12]. Part 2 is an easy consequence, see below.
Lemma C.1.
Let $f:\mathbb{D}\to\mathbb{D}$ be univalent. For some $\zeta\in S^{1}$, suppose that the angular limit $f(\zeta)$ exists and $f(\zeta)\in S^{1}$. Assume that the angular derivative $f^{\prime}(\zeta)$ is finite.
  1. (b1)
    If $f(0)=0$, then
    (C.1) $|f^{\prime}(0)|^{-1/2}\leq|f^{\prime}(\zeta)|.$
  2. (b2)
    in general,
    (C.2) $\frac{(1-|f(0)|)^{3/2}}{(1+|f(0)|)^{1/2}}|f^{\prime}(0)|^{-1/2}\leq|f^{\prime}(\zeta)|.$
Proof of (b2): let $g(z)=\frac{f(z)-A}{1-\bar{A}f(z)}$ where $A=f(0)$. As $g^{\prime}(z)=\frac{f^{\prime}(z)(1-|A|^{2})}{(1-\bar{A}f(z))^{2}}$, then $g^{\prime}(0)=\frac{f^{\prime}(0)}{1-|A|^{2}}$ and $|g^{\prime}(\zeta)|=\frac{|f^{\prime}(\zeta)|(1-|A|^{2})}{|1-\bar{A}f(\zeta)|^{2}}\leq\frac{|f^{\prime}(\zeta)|(1-|A|^{2})}{(1-|A|)^{2}}=\frac{|f^{\prime}(\zeta)|(1+|A|)}{(1-|A|)}$. By (C.1), this implies (C.2).
The following lemma and its corollary state that, if $E,f(E)\subset S^{1}$, then the size of $E$ is small provided $|f^{\prime}(0)|$ is small and $f(0)$ is away from the unit circle. Exact bounds are obtained in [15], under the assumption that $E$ is contained in an open set $E_{0}\subset S^{1}$ such that $f(E_{0})\subset S^{1}$ (and $f(0)=0$).
Lemma C.2.
Let $f:\mathbb{D}\to\mathbb{D}$ be univalent. Let $E\subset S^{1}$ be a measurable set such that $f(E)\subset S^{1}$, and $|E|>0$. Then for every $\epsilon>0$ there exists a closed set $L\subset E$ such that $|L|>|E|-\epsilon$, $f(L)$ is measurable and
(C.3) $|f(L)|\geq|L|\sqrt{\frac{(1-|f(0)|)^{3}}{(1+|f(0)|)|f^{\prime}(0)|}}.$
Proof.
As $f^{\prime}(\zeta)$ exists and is finite for a.e. $\zeta\in E$, and $f^{\prime}(\zeta)=\lim_{n\to\infty}f^{\prime}((1-1/n)\zeta)$ where $f^{\prime}$ is continuous in $\mathbb{D}$, then $f^{\prime}:E\to\mathbb{C}$ is measurable. By Lusin’s theorem, there exists a closed subset $E^{\prime}$ of $E$ such that $|E^{\prime}|>|E|-\epsilon/2$ and $f^{\prime}:E^{\prime}\to\mathbb{C}$ is (uniformly) continuous. Fix an angle $\beta\in(0,\pi/2)$ and a radius $r\in(0,\sin\beta)$.
Given $X\subset S^{1}$, let
$W_{X}=\cup_{a\in X}\Delta_{a}(\beta,r)$
(see Definition B.1)). For every $n$ big enough, define $M_{n}:E^{\prime}\to\mathbb{R}$ as follows:
$M_{n}(a)=\sup\{|f^{\prime}(z)-f^{\prime}(a)|:z\in\Delta_{a}(\beta,1-1/n).$
Since $f^{\prime}(z)\to f^{\prime}(a)$ as $z\to a$ while $|\arg(z-a)|<\beta$, $M_{n}(a)\to 0$ for every $a\in E^{\prime}$. By Egorov’s theorem, there exists a closed set $L\subset L^{\prime}$, $|L|>|L^{\prime}|-\epsilon/2$, such that $M_{n}\to 0$ uniformly on $L$. (Note that $|E|>|E|-\epsilon$.) Consider $f^{\prime}$ on $W_{L}$. We show that $f^{\prime}(z)\to f^{\prime}(a)$ as $z\to a$ in $W_{L}$ uniformly in $a\in L$. Indeed, given a small $t>0$, choose a small $\delta_{0}>0$ such that $|f^{\prime}(b)-f^{\prime}(a)|<t$ whenever $a,b\in L$ and $|b-a|<\delta_{0}$, and choose $N$ such that $M_{N}(b)<t/2$ for all $b\in L$. Now, let $\delta=\min\{1/N,\delta_{0}/(2\cos\beta)\}$. It is chosen so that, for any $z\in B(1,\delta)\cap\mathbb{D}$, if $z\in\Delta_{c}(\beta,r)$ for some $c\in S^{1}$ then $|c-1|<\delta_{0}$.
If now $|z-a|<\delta$ for some $a\in L$, $z\in W_{L}$, then, by the definition of $W_{L}$, $z\in\Delta_{b}(\beta,r)$ for some $b\in L$, on the other hand, $|b-a|<\delta_{0}$, by the choice of $\delta$. Therefore,
$|f^{\prime}(z)-f^{\prime}(a)|\leq|f^{\prime}(z)-f^{\prime}(b)|+|f^{\prime}(b)-f^{\prime}(a)|<M_{N}(b)+t/2<t/2+t/2=t.$
Thus $f^{\prime}(z)\to f^{\prime}(a)$ as $z\to a$ for $z\in W_{L}$, uniformly in $a\in L$. Finally, as $f^{\prime}$ is continuous in $\mathbb{D}$, we conclude that $f^{\prime}$ is a continuous (in particular, bounded) function in the closure $\overline{W_{L}}$ of $W_{L}$.
Since $L=\overline{W_{L}}\cap S^{1}$, it is enough to prove (C.3) with $L$ replaced by $F\subset L$ such that $W_{F}$ is a component of the (open) set $W_{E}$. Fix such an $F$. $W_{F}$ is a simply connected domain, hence, there exists a conformal isomorphism $R:\mathbb{D}\to W_{F}$. As $W_{F}$ is bounded by a rectifiable Jordan curve, by Riesz-Privalov’s theorem, the derivative $R^{\prime}$ belongs in the Hardy space $H^{1}(\mathbb{D})$. Hence, $R^{-1}(F)$ is measurable. Consider a univalent map $g:=f\circ R:\mathbb{D}\to\mathbb{D}$. As $f^{\prime}$ is bounded in $W_{F}$, $g^{\prime}\in H^{1}(\mathbb{D})$ as well and the set $g(R^{-1}(F))$ is measurable. By Riesz-Privalov’s theorem, we can write:
$|f(F)|=|g(R^{-1}(F))|=\int_{R^{-1}(F)}|g^{\prime}(w)||dw|=\int_{R^{-1}(F)}|f^{\prime}(R(w))||R^{\prime}(w)||dw|=$
$\int_{F}|f^{\prime}(z)||dz|\geq|F|\sqrt{\frac{(1-|f(0)|)^{3}}{(1+|f(0)|)|f^{\prime}(0)|}},$
where the latter inequality follows from Lemma C.1.
Corollary C.3.
Let $f:\mathbb{D}\to\mathbb{D}$ be univalent. Let $E\subset S^{1}$ be a measurable set such that $f(E)\subset S^{1}$, and $|E|>0$. Then
(C.4) $|E|\sqrt{\frac{(1-|f(0)|)^{3}}{(1+|f(0)|)|f^{\prime}(0)|}}\leq 2\pi.$